# “Feeling that I was Collaborating with Them:” A 20-year Systematic Literature Review of Social Virtual Reality Leveraging Collaboration

### v2

@@ Abstract.
As more people meet, interact, and socialize online, Social Virtual Reality (VR) emerges as a promising technology that can bridge the gap between traditional face-to-face and online communication. Compared to traditional screen-based applications, Social VR provides immersive, spatial, and three-dimensional social interactions, making it a promising tool for enhancing collaborations. To map the existing research in this domain, we conducted a 20-year systematic literature review to characterize how Social VR has been employed for collaboration. After screening 2,035 articles, we identified 62 articles that addressed how Social VR has supported collaboration among remote users. Our findings show that Social VR can enhance team collaboration on three key levels: enhancing individual perceptions and experiences within their groups, fostering team dynamics with virtual elements that enable realistic interactions, and employing affordances unique in VR that augment users’ spaces. Future research should explore how Social VR can support long-term collaboration, foster trust, enable more diverse and inclusive participation, and move beyond replicating physical-world interactions by leveraging the unique affordances of immersive environments. This review highlights the current practices, challenges, and future research opportunities within CSCW, offering insights for theorizing the impact of Social VR on team collaboration and for designing new applications that effectively support remote collaborations.
@@
copyright: acmlicensedjournal: PACMHCIjournalyear: 2025journalvolume: 4journalnumber: CSCW2publicationmonth: 10price: 15.00ccs: Human-centered computing Collaborative and social computingccs: Human-centered computing Human computer interaction (HCI)
[…]
Figure 1. Overview of the systematic literature review findings on Social Virtual Reality (VR) collaboration. The figure illustrates the three main themes: (1) Individual Perception and Experience of Collaboration in Social VR, examining sense of presence and avatar perception; (2) Team Dynamics and Collaboration in Virtual Environments, focusing on social presence, collaboration realness, trust-building, and engagement; and (3) Enabling Collaboration through Social VR Affordances, addressing virtual environment functionality, virtual objects and tool facilitation, and immersion enhancement through spatial audio and movement.
@@ 1. Introduction
The rise of Social Virtual Reality (Social VR) is reshaping how people can communicate and collaborate in digital spaces. In February 2022, Meta’s social VR Horizon Worlds achieved 300,000 active monthly users in the US (Heath, 2022) and the platform Rec Room surpassed 3 million monthly active users (Feltham, 2022), showing the growing adoption of Social VR as an environment for social connections.
@@
Unlike traditional screen-based communication tools, such as video conferencing or instant messaging, Social VR enables users to immerse themselves in three-dimensional spaces through head-mounted displays (HMDs), creating a sense of physical presence and embodied interaction in simulated environments (Maloney and Freeman, 2020). Social VR offers unique affordances that can enhance communication among distributed groups, such as avatar embodiment, shared object manipulation, spatial audio, and interactive environments (Li et al., 2021; Freeman et al., 2022). Its capacity to simulate real-world experiences has positioned Social VR as a transformative medium for social interactions, particularly after the COVID-19 pandemic, which increased the need for robust alternatives to face-to-face interactions (Waizenegger et al., 2020; Ball, 2022; Gómez-Zará et al., 2023).
@@
Despite the growing interest in Social VR, its use in collaborative and task-oriented settings remains unclear. While much of the existing research has focused on its applications in gaming, training, and casual gatherings (Bujić et al., 2021; Angel-Urdinola et al., 2021; Maloney and Freeman, 2020), the potential of Social VR for enabling collaborative meetings and teamwork has been underexplored. Companies such as Microsoft have attempted to integrate Social VR environments into professional settings to support meetings, conversations, and teamwork (Tilley, 2022; Roth, 2023). However, adoption has often fallen short due to several challenges, including low-quality simulated environments, steep learning curves, cognitive dissonance between physical and virtual worlds, and a lack of applications specifically designed for professional purposes (Ashtari et al., 2020; Zhao, 2011; Hubbard and Aguinis, 2023). Moreover, the literature has yet to fully explore how these immersive technologies support team processes, such as trust-building, long-term collaboration, and group cohesion (Maloney and Freeman, 2020; Abramczuk et al., 2023; Osborne et al., 2023). These challenges highlight important gaps in both the design and understanding of Social VR as an environment for distributed collaboration. Specifically, there is a need to better characterize how Social VR environments can address these limitations and replicate the richness of in-person interactions in group settings.
@@
Characterizing studies on Social VR that address collaboration and teamwork is important to understanding its potential as a transformative tool for distributed collaboration. Research communities in Computer-Supported Cooperative Work (CSCW) and Human-Computer Interaction (HCI) have long sought to design and study systems that support collaboration among remote groups (Olson and Olson, 2000; Harris et al., 2019). As such, Social VR offers an opportunity to rethink how digital tools can enhance collaboration in remote and hybrid contexts, bridging the limitations of current systems while using its unique affordances for presence and engagement. This paper provides a comprehensive overview of the current research landscape in Social VR, identifying critical gaps in areas such as trust-building, user engagement, scalability of teamwork, long-term collaboration, the diversity of study samples, and the design of effective applications to support various stages of collaborative processes. By synthesizing these findings, we aim to highlight actionable opportunities for future CSCW research and offer design implications for developing effective collaborative VR systems that address the limitations of existing tools. This effort is intended to bridge the gap between theoretical advancements and practical applications, helping organizations and researchers leverage Social VR for more meaningful and productive remote teamwork. Given this motivation, the research questions that guide this study are:
@@
To address these research questions, we conducted a systematic literature review on Social VR addressing aspects of collaboration. Our goal is to identify key solutions, insights, and recommendations that support group interaction and teamwork in Social VR contexts. This systematic review covered articles published in the IEEE Xplore Digital Library, ACM Digital Library, and Web of Science databases from 2004 to 2024, covering 20 years of research. After screening 2,035 articles, we identified 62 articles that specifically focused on how Social VR has supported teams and enhanced collaboration. Based on this corpus, we conducted a thematic analysis and identified three core themes: (1) “Individual Perception and Experience of Collaboration in Social VR,” (2) “Team Dynamics and Collaboration in Virtual Environments,” and (3) “Enabling Collaboration through Social VR Affordances.” These themes illustrate how Social VR’s affordances and mechanisms at the individual, team, and environment levels can facilitate collaboration.
Our systematic review also highlights several research gaps, including insufficient attention to diverse user populations across demographics and technical expertise, limited understanding of how Social VR scales to larger teams and complex tasks, and a lack of standardized evaluation frameworks for assessing collaboration. Moreover, the themes uncover key questions about trust-building mechanisms, team dynamics, and the transition from traditional online platforms to Social VR for collaboration purposes. Overall, this review provides an comprehensive assessment of the current capabilities and promising directions that Social VR can offer to people and organizations in real-world collaborative settings.
The contributions of this paper are twofold. First, it offers a comprehensive review of the use of Social VR environments to support remote collaboration, encompassing an analysis of applications and their impact on teamwork. Second, it provides a holistic map for future CSCW research and system development in Social VR to enhance collaboration. The themes and implications discussed in this study aim to inform CSCW designers and researchers about the opportunities and challenges involved in making Social VR effective for remote collaboration.
@@ 2.1. Teams and Teamwork Technologies
Teams are groups of two or more individuals working interdependently, adaptively, and dynamically toward a common and valued goal, bringing diverse skills, knowledge, and perspectives to address challenging tasks (Gómez-Zará et al., 2020; Salas et al., 2000). Unlike individuals working alone, teams can bring together people with diverse skills and knowledge to achieve better outcomes (Engel et al., 2015; Mathieu et al., 2017). Team members work on different parts of a task, relying on each other, sharing responsibility to achieve a common goal, and resolving complex problems in a distributed manner (Mathieu et al., 2017; Gómez-Zará et al., 2020; Salas et al., 2000). Previous research shows that successful teamwork relies on clear communication, coordination, mutual support, and individual contributions aligned with the overall task strategy (Mathieu et al., 2017). While teamwork requires more coordination and collaboration than individual efforts, it provides a higher capacity to manage challenging tasks, particularly when specialized skills are needed (Mathieu et al., 2017). The importance of teams has grown as organizational structures have shifted over the past decades from individual roles to team-based systems (Wuchty et al., 2007), with tasks increasingly designed around groups working toward shared objectives (Arora et al., 2023; Mathieu et al., 2017). For instance, in IT organizations, project teams require programmers, designers, and project managers to work interdependently on complex software development projects, where effective communication, task coordination, and shared understanding are essential for delivering successful outcomes (Arora et al., 2023; Huang et al., 2023).
@@
The impact of online technologies on collaboration and teamwork continues to evolve alongside the dynamic technological ecosystem (Dulebohn and Hoch, 2017; Jimenez et al., 2017; Treem et al., 2020). Advancements in groupware, intranets, social platforms, smartphones, cloud technologies, and videoconferencing platforms have redefined the degrees of “virtualness” (Staples and Webster, 2008; Brown et al., 2020) in which teams operate, ranging from purely face-to-face to only virtual interactions (Griffith and Neale, 2001; Brown et al., 2020). For example, hybrid workplaces allow team members to work remotely and meet in person on specific occasions (Brown et al., 2020). Workers can still be physically in one building, but coordinate through online communication. Since team members do not necessarily work in the same physical place, current challenges that teams face reside in coordinating, staying cohesive, and performing effectively (Gilson et al., 2015).
@@
With the rise of hybrid and remote work (Freeman et al., 2022; Bjørn et al., 2014; Breideband et al., 2023), online collaboration tools are now key components for facilitating effective communication and teamwork (Cai and Tanaka, 2019; Li et al., 2021). Instant messaging platforms, like Microsoft Teams and Slack, provide another layer of group communication by enabling quick exchanges of information and facilitating asynchronous collaboration (Cai and Tanaka, 2019; Li et al., 2021). These online communication platforms support various forms of media sharing and integrate with other productivity tools, enhancing workflow efficiency and group awareness. However, they may contribute to information overload and can sometimes lead to misinterpretations without the contextual cues present in verbal communication (Cha et al., 2018). Video conferencing platforms (e.g., Zoom, Google Meets, Microsoft Teams) also create a shared visual space that enhances user engagement and presence, facilitating real-time communication from different locations (Tang et al., 2023; Karis et al., 2016). Despite offering significant benefits, they often lack the depth of non-verbal cues available in face-to-face interactions, such as body language (Freeman and Acena, 2021; Bente et al., 2008; Kanawattanachai and Yoo, 2002), and can increase the mental exhaustion and cognitive load for their prolonged use (Bailenson, 2021). Previous research has also shown that remote collaboration can curb several activities that require group thinking, such as creativity (Brucks and Levav, 2022), idea generation (Lin et al., 2023), identity (Bartel et al., 2007), and cohesion (Yang et al., 2022). These current limitations have motivated CSCW researchers to explore how immersive environments, such as Social VR, can improve remote collaboration and enhance the quality of remote teamwork.
@@ 2.2. Social VR
Social Virtual Reality (VR) refers “to 3D virtual spaces where multiple users can interact with each other using VR head-mounted displays (HMDs)” (Maloney and Freeman, 2020). Social VR encompasses a subset of VR platforms and applications specifically designed to support social interaction and co-presence (Li et al., 2021; Freeman and Maloney, 2021), where users can feel as if they are physically present with others in the same virtual space (McVeigh-Schultz et al., 2019). Given these unique affordances, Social VR has the potential to change how people engage, communicate, and collaborate in various activities, from social gatherings and entertainment to educational and professional teamwork settings (Deighan et al., 2023). Platforms like ‘VRChat’ have become popular spaces for individuals to connect in VR environments and foster a sense of community and well-being, particularly with the emergence of the COVID-19 pandemic (Osborne et al., 2023).
@@
Unlike traditional video conferencing applications, Social VR allows people to immerse themselves in simulated environments to interact and socialize in more interactive ways (Freeman and Acena, 2021). Social VR applications can enable remote users to interact with 3D virtual objects in immersive spaces (Freeman et al., 2022; Li et al., 2021), utilize avatars that can differ from their real bodies and appearances, employ interactive tools that break the physical laws (Freeman et al., 2022), and help people stay more engaged in the team conversation using their bodies and movements (Abramczuk et al., 2023; Olaosebikan et al., 2022; Steinicke et al., 2020). The unique affordances that VR provides can help users communicate not only with words but also with gestures, body language, and eye gaze with the help of body tracking technologies (Freeman et al., 2022).
@@
Previous research shows that Social VR can enable people to interact in similar ways to face-to-face settings, helping people build stronger social bonds (Yoon and Leem, 2021). For entertainment purposes, VR has demonstrated its potential to enhance shared experiences. For example, Lee et al. (Lee et al., 2022) developed a VR lobby for digital opera social experiences, highlighting how Social VR can replicate and enrich live entertainment through immersive environments. Beyond entertainment, Young et al. (Young et al., 2023) demonstrated that Social VR can facilitate remote learning by enhancing students’ participation in online discussions and activities, improving their sense of presence that traditional online learning platforms often do not provide. In another study, Sanaei et al. (Sanaei et al., 2023) conducted a controlled experiment to compare groups working on a circuit repair in face-to-face (F2F), video conferencing, and VR. The study found that participants in VR had a significantly better task experience and lower cognitive efforts compared to participants in the video conferencing condition. These studies suggest that VR can emulate the interactive aspects of in-person interactions that traditional online communication platforms lack.

### v3

# “Feeling that I was Collaborating with Them:” A 20-year Scoping Review of Social Virtual Reality Leveraging Collaboration

@@ Abstract.
As more people meet, interact, and socialize online, Social Virtual Reality (VR) emerges as a technology that bridges the gap between traditional face-to-face and online communication. Unlike traditional screen-based applications, Social VR provides immersive, spatial, and three-dimensional social interactions, making it a potential tool for enhancing remote collaborations. Despite the growing interest in Social VR, research on its role in collaboration remains fragmented, calling for a synthesis to identify research gaps and future directions. We conducted a 20-year scoping review, screening 2,035 articles and identifying 62 articles that addressed how Social VR has supported collaboration. Our analysis shows three key levels of support: Social VR can enhance individual perceptions and experiences within their groups, foster team dynamics with virtual elements that enable realistic interactions, and employ the unique affordances of VR to augment users’ spaces. We discuss how future research in Social VR should move beyond replicating physical-world interactions and explore how immersive environments can cultivate long-term collaboration, trust, and more diverse and inclusive participation. This review highlights the current practices and challenges, highlighting new opportunities for theorizing and designing Social VR systems that responsibly support remote collaborations.
@@
copyright: ccjournal: PACMHCIjournalyear: 2026journalvolume: 10journalnumber: 2article: CSCW017publicationmonth: 4doi: 10.1145/3788053ccs: Human-centered computing Collaborative and social computingccs: Human-centered computing Human computer interaction (HCI)
Refer to caption Illustrated overview of a scoping review on Social Virtual Reality collaboration. The figure shows three illustrated themes: individual perception and experience in Social VR, team dynamics and collaborative processes in virtual environments, and enabling collaboration through tools and affordances, depicted through people interacting in immersive virtual settings.
Figure 1. Overview of the scoping review findings on Social Virtual Reality (VR) collaboration. The figure illustrates the three main themes: (1) Individual Perception and Experience of Collaboration in Social VR, examining sense of presence and avatar perception; (2) Team Dynamics and Collaboration in Virtual Environments, focusing on social presence, collaboration realness, trust-building, and engagement; and (3) Enabling Collaboration through Social VR Affordances, addressing virtual environment functionality, virtual objects and tool facilitation, and Supporting Presence through spatial audio and movement.
@@ 1. Introduction
The rise of Social Virtual Reality (Social VR) is transforming the way people communicate and collaborate in digital spaces. In February 2022, Meta’s social VR Horizon Worlds achieved 300,000 active monthly users in the US (Heath, 2022) and the platform ‘Rec Room’ surpassed 3 million monthly active users (Feltham, 2022), showing the growing adoption of Social VR as an environment for social connections. Unlike traditional screen-based communication tools—such as video conferencing or instant messaging—Social VR enables users to immerse themselves in three-dimensional spaces through head-mounted displays (HMDs), creating a sense of physical presence and embodied interaction in simulated environments (Maloney and Freeman, 2020). It offers unique affordances that can enhance communication among distributed groups, such as avatar embodiment, shared object manipulation, spatial audio, and interactive environments (Li et al., 2021; Freeman et al., 2022). Its capacity to simulate real-world experiences has positioned Social VR as a transformative medium for social interactions, particularly after the COVID-19 pandemic, which increased the need for robust alternatives to face-to-face interactions (Waizenegger et al., 2020; Ball, 2022; Gómez-Zará et al., 2023).
@@
Despite the growing interest in Social VR, its use in collaborative and task-oriented settings remains unclear. Although most existing research has focused on its applications in gaming, training, and casual gatherings (Bujić et al., 2021; Angel-Urdinola et al., 2021; Maloney and Freeman, 2020), the potential of Social VR to enable collaborative meetings and teamwork has been underexplored. Companies like Microsoft have attempted to integrate Social VR environments into professional settings to support meetings, conversations, and teamwork (Tilley, 2022; Roth, 2023). However, adoption has often fallen short due to several challenges, including low-quality simulated environments, steep learning curves, cognitive dissonance between physical and virtual worlds, and a lack of applications specifically designed for professional purposes (Ashtari et al., 2020; Zhao, 2011; Hubbard and Aguinis, 2023). Moreover, the literature has not yet fully explored how these technologies support team processes, such as trust-building, long-term collaboration, and group cohesion (Maloney and Freeman, 2020; Abramczuk et al., 2023; Osborne et al., 2023). These challenges highlight significant gaps in both the design and understanding of Social VR as an environment for distributed collaboration. Specifically, there is a need to better characterize how Social VR environments can address these limitations and replicate the richness of in-person interactions in group settings.
@@
Characterizing studies on Social VR’s collaborative uses is crucial for understanding its potential as a transformative tool for distributed collaboration. Research communities in Computer-Supported Cooperative Work (CSCW) and Human-Computer Interaction (HCI) have long sought to design and study systems that support collaboration among remote groups (Olson and Olson, 2000; Harris et al., 2019). As such, Social VR offers an opportunity to rethink how digital tools can enhance collaboration in remote and hybrid contexts, bridging the limitations of current systems while using its unique affordances for presence and engagement. This paper provides a comprehensive overview of the current research landscape in Social VR, identifying critical gaps in areas such as trust-building, user engagement, teamwork scalability, long-term collaboration, the diversity of study samples, and the design of effective applications to support various stages of collaborative processes. By synthesizing these findings, we aim to highlight actionable opportunities for future CSCW research and offer design implications for developing effective collaborative VR systems that address the limitations of existing tools. Given this motivation, the research questions that guide this study are:
@@
To address these research questions, we conducted a scoping review, a methodology designed to map the breadth and diversity of existing research. Unlike systematic reviews, which assess evidence for narrowly defined research questions, our scoping review aims to characterize the landscape and identify solutions and insights that support teamwork in Social VR (Munn et al., 2018). This review covered articles published in the IEEE Xplore Digital Library, ACM Digital Library, and Web of Science databases from 2004 to 2024, covering 20 years of research. After screening 2,035 articles, we identified 62 articles that specifically focused on how Social VR has supported teams and enhanced collaboration. Based on this corpus, we conducted a thematic analysis and identified three core themes: (1) “Individual Perception and Experience of Collaboration in Social VR,” (2) “Team Dynamics and Collaboration in Virtual Environments,” and (3) “Enabling Collaboration through Social VR Affordances.” These themes illustrate how Social VR’s affordances and mechanisms at the individual, team, and environment levels can facilitate collaboration.
Our scoping review also highlights several research gaps, including insufficient attention to diverse user populations across demographics and technical expertise, limited understanding of how Social VR scales to larger teams and complex tasks, and a lack of standardized evaluation frameworks for assessing collaboration. Moreover, the themes uncover key questions about trust-building mechanisms, team dynamics, and the transition from traditional online platforms to Social VR for collaboration purposes. Overall, this review provides a comprehensive assessment of the current capabilities and directions that Social VR can offer to individuals and organizations in real-world collaborative settings.
The contributions of this paper are twofold. First, it provides a comprehensive review of how Social VR has been designed and used to support remote collaboration, analyzing both its applications and its impact on teamwork. Second, it provides a holistic road map for advancing Social VR research and system development to strengthen collaboration. Together, the themes and implications discussed in this study provide guidance for CSCW researchers and practitioners seeking to harness the opportunities and address the challenges of making Social VR effective for remote collaboration.
@@ 2.1. Teams and Teamwork Technologies
Teams are groups of two or more individuals working interdependently, adaptively, and dynamically towards a common and valued goal, bringing diverse skills, knowledge, and perspectives to address challenging tasks (Gómez-Zará et al., 2020; Salas et al., 2000). Unlike individuals working alone, teams can bring together people with diverse skills and knowledge to achieve better outcomes (Engel et al., 2015; Mathieu et al., 2017). Team members work on different parts of a task, relying on each other, sharing responsibility to achieve a common goal, and resolving complex problems in a distributed manner (Mathieu et al., 2017; Gómez-Zará et al., 2020; Salas et al., 2000). Previous research shows that successful teamwork relies on clear communication, coordination, mutual support, and individual contributions aligned with the overall task strategy (Mathieu et al., 2017). While teamwork requires more coordination and collaboration than individual efforts, it provides a higher capacity to manage challenging tasks, particularly when specialized skills are needed (Mathieu et al., 2017). The importance of teams has grown as organizational structures have shifted over the past decades from individual roles to team-based systems (Wuchty et al., 2007), with tasks increasingly designed around groups working toward shared objectives (Arora et al., 2023; Mathieu et al., 2017). For instance, in IT organizations, project teams require programmers, designers, and project managers to work interdependently on complex software development projects, where effective communication, task coordination, and shared understanding are essential for delivering successful outcomes (Arora et al., 2023; Huang et al., 2023).
@@
The impact of online technologies on collaboration and teamwork continues to evolve alongside the dynamic technological ecosystem (Dulebohn and Hoch, 2017; Jimenez et al., 2017; Treem et al., 2020). Advancements in groupware, intranets, social platforms, smartphones, cloud technologies, and videoconferencing platforms have redefined the degrees of “virtualness” (Staples and Webster, 2008; Brown et al., 2020) in which teams operate, ranging from purely face-to-face to only virtual interactions (Griffith and Neale, 2001; Brown et al., 2020). For example, hybrid workplaces allow team members to work remotely and meet in person on specific occasions (Brown et al., 2020). Workers can still be physically present in one building, but coordinate through online communication. Since team members do not necessarily work in the same physical place, current challenges that teams face include coordinating, staying cohesive, and performing effectively (Gilson et al., 2015).
@@
With the rise of hybrid and remote work (Freeman et al., 2022; Bjørn et al., 2014; Breideband et al., 2023), online collaboration tools are now key components in facilitating effective communication and teamwork (Cai and Tanaka, 2019; Li et al., 2021). Instant messaging platforms, such as Microsoft Teams and Slack, provide another layer of group communication by enabling quick exchanges of information and facilitating asynchronous collaboration (Cai and Tanaka, 2019; Li et al., 2021). These online communication platforms support various forms of media sharing and integrate with other productivity tools, enhancing workflow efficiency and group awareness. However, they may contribute to information overload and can sometimes lead to misinterpretations without the contextual cues present in verbal communication (Cha et al., 2018). Video conferencing platforms (e.g., Zoom, Google Meets, Microsoft Teams) also create a shared visual space that enhances user engagement and presence, facilitating real-time communication from different locations (Tang et al., 2023; Karis et al., 2016). Despite offering significant benefits, they often lack the depth of non-verbal cues available in face-to-face interactions, such as body language (Freeman and Acena, 2021; Bente et al., 2008; Kanawattanachai and Yoo, 2002), and can increase mental exhaustion and cognitive load with prolonged use (Bailenson, 2021). Previous research has also shown that remote collaboration can curb several activities that require group thinking, such as creativity (Brucks and Levav, 2022), idea generation (Lin et al., 2023), identity (Bartel et al., 2007), and cohesion (Yang et al., 2022a). These current limitations have motivated CSCW researchers to explore how immersive environments, such as Social VR, can improve remote collaboration and enhance the quality of remote teamwork.
@@ 2.2. Social VR
Social Virtual Reality (VR) refers to “3D virtual spaces where multiple users can interact with each other using VR head-mounted displays (HMDs)” (Maloney and Freeman, 2020). Social VR encompasses a subset of VR platforms and applications specifically designed to support social interaction and co-presence (Li et al., 2021; Freeman and Maloney, 2021), where users can feel as if they are physically present with others in the same virtual space (McVeigh-Schultz et al., 2019). Given these unique affordances, Social VR has the potential to change how people engage, communicate, and collaborate in various activities, from social gatherings and entertainment to educational and professional teamwork settings (Deighan et al., 2023). Platforms like ‘VRChat’ have become popular spaces for individuals to connect in VR environments and foster a sense of community and well-being, particularly with the emergence of the COVID-19 pandemic (Osborne et al., 2023).
@@
Unlike traditional video conferencing applications, Social VR enables people to immerse themselves in simulated environments, allowing for more interactive ways to interact and socialize (Freeman and Acena, 2021). Social VR applications can enable remote users to interact with 3D virtual objects in immersive spaces (Freeman et al., 2022; Li et al., 2021), utilize avatars that can differ from their real bodies and appearances, employ interactive tools that break the physical laws (Freeman et al., 2022), and help people stay more engaged in the team conversation using their bodies and movements (Abramczuk et al., 2023; Olaosebikan et al., 2022; Steinicke et al., 2020). The unique affordances that VR provides can help users communicate not only with words but also with gestures, body language, and eye gaze with the help of body tracking technologies (Freeman et al., 2022).
@@
Previous research shows that Social VR can enable people to interact in similar ways to face-to-face settings, helping people build stronger social bonds (Yoon and Leem, 2021). For entertainment purposes, VR has demonstrated its potential to enhance shared experiences. For example, Lee et al. (Lee et al., 2022) developed a VR lobby for digital opera social experiences, highlighting how Social VR can replicate and enrich live entertainment through immersive environments. Beyond entertainment, Young et al. (Young et al., 2023) demonstrated that Social VR can facilitate remote learning by enhancing students’ participation in online discussions and activities, improving their sense of presence that traditional online learning platforms often do not provide.
In another study, Sanaei et al. (Sanaei et al., 2023) conducted a controlled experiment to compare groups working on a circuit repair in face-to-face (F2F), video conferencing, and VR. The study found that participants in VR had a significantly better task experience and lower cognitive efforts compared to participants in the video conferencing condition. These studies suggest that VR can emulate the interactive aspects of in-person interactions that traditional online communication platforms lack.